# A Retrospective Systematic Study on Hierarchical Sparse Query Transformer-assisted Ultrasound Screening for Early Hepatocellular Carcinoma

### v1

@@ Abstract
Hepatocellular carcinoma (HCC) ranks as the third leading cause of cancer-related mortality worldwide, with early detection being crucial for improving patient survival rates. However, early screening for HCC using ultrasound suffers from insufficient sensitivity and is highly dependent on the expertise of radiologists for interpretation. Leveraging the latest advancements in artificial intelligence (AI) in medical imaging, this study proposes an innovative Hierarchical Sparse Query Transformer (HSQformer) model that combines the strengths of Convolutional Neural Networks (CNNs) and Vision Transformers (ViTs) to enhance the accuracy of HCC diagnosis in ultrasound screening. The HSQformer leverages sparse latent space representations to capture hierarchical details at various granularities without the need for complex adjustments, and adopts a modular, plug-and-play design philosophy, ensuring the model’s versatility and ease of use. The HSQformer’s performance was rigorously tested across three distinct clinical scenarios: single-center, multi-center, and high-risk patient testing. In each of these settings, it consistently outperformed existing state-of-the-art models, such as ConvNext and SwinTransformer. Notably, the HSQformer even matched the diagnostic capabilities of senior radiologists and comprehensively surpassed those of junior radiologists. The experimental results from this study strongly demonstrate the effectiveness and clinical potential of AI-assisted tools in HCC screening. The full code is available at https://github.com/Asunatan/HSQformer.
footnotetext: # Equal contribution.footnotetext: * Corresponding Author.
@@ 1 Introduction
Liver cancer is the sixth most common cancer worldwide and the third leading cause of cancer-related deaths, with hepatocellular carcinoma (HCC) being the most prevalent type of primary liver cancer[1]. Early detection and treatment of HCC can improve patient survival rates and life expectancy. Ultrasound (B-mode) is the preferred imaging modality for screening high-risk populations for HCC. Compared to CT or MRI, ultrasound offers the advantages of portability, radiation-free imaging, and real-time visualization. However, according to meta-analyses, the sensitivity of conventional B-mode ultrasound in detecting HCC ranges from only 59% to 78%, which is highly dependent on the clinical experience of radiologists[2, 3]. Moreover, ultrasound examinations require substantial time from these specialists for image interpretation and secondary confirmation, often making clinical assessments inefficient. In recent years, with the rapid development of AI, deep learning-based computer-aided diagnosis (CAD) systems have shown great potential. These systems can help improve scanning efficiency and reduce the dependency of diagnostic outcomes on the clinical experience of radiologists or sonographers.
@@
Existing computer-aided diagnosis systems are primarily based on Convolutional Neural Networks (CNNs) or Vision Transformers (ViTs)[4]. CNNs have been pivotal in the field of computer vision, particularly for tasks involving image classification, object detection and semantic segmentation, due to their proficiency in capturing local texture information. However, their performance is often constrained by the size of their receptive fields, which limits their ability to perceive long-range dependencies and integrate global contextual information. To overcome these limitations, researchers have explored various techniques, such as enlarging convolutional kernels[5, 6, 7], employing structural re-parameterization[8, 9, 10], and utilizing dilated or deformable convolutions[11, 12, 13]. In contrast, ViTs have emerged as a powerful alternative, particularly excelling in capturing global semantic information. The self-attention mechanism inherent in ViTs facilitates a comprehensive understanding of input data by evaluating the interactions between all elements. Nevertheless, ViTs often lack the inductive bias that CNNs possess, which is crucial for learning from limited data[14, 15]. As a result, ViTs typically require extensive datasets and significant computational resources for effective pre-training. To mitigate computational demands while maintaining performance, windowed attention mechanisms[16, 17, 18] and pooling operations[19, 20, 21] have been proposed. This raises the question: is there a straightforward method to address the limitations of these two architectures? A simple yet feasible solution is to combine the strengths of CNNs and ViTs to form a hybrid model.
@@
The aforementioned methods typically require modifications to the network architecture or the use of intricate design techniques. Is it possible to achieve effective complementarity between CNNs and ViTs without altering the network structure? This task presents significant challenges. As shown in Figure 1, the two most popular network architectures, ConvNext (CNNs)[22] and SwinTransformer (ViTs)[16], were selected for preliminary experiments. We observed that the benefits of simply combining[23] CNNs and ViTs were minimal. Our experience suggests that late fusion tends to focus excessively on high-level semantics relevant to classification tasks, while neglecting low-level details and textures. Moreover, this approach fails to leverage multi-scale features, which have been proven crucial for downstream tasks such as segmentation and detection. A staged fusion approach might seem intuitive; however, previous studies have shown that combining CNNs and ViTs in this way can lead to feature redundancy[24, 23].
Inspired by latent space feature representation and sparse learning, we propose an innovative Hierarchical Sparse Query Transformer (HSQformer), designed to effectively integrate CNNs and ViTs for enhancing real-world clinical ultrasound HCC screening. The HSQformer is built on two backbone feature extractors, ConvNeXt and Swin Transformer, and adopts a modular, plug-and-play approach. This design requires no additional bells and whistles, thereby ensuring ease of extensibility. Specifically, the HSQformer stacks modules of Cross-attention, Self-attention, and Mixture of experts (CSM) to explore AI’s potential in aiding diagnosis. Within this framework, the cross-attention mechanism facilitates inter-feature communication, the self-attention module enhances intra-feature representations, and the mixture of experts specializes in different aspects of the data through the incorporation of sparse learning.
To validate the real-world effectiveness of HSQformer, we retrospectively collected data from 2014 to 2022 and constructed three clinical scenarios: single-center, multi-center, and high-risk patient testing. Our proposed model achieves state-of-the-art performance across all three datasets, outperforming previous best methods such as ConvNext and SwinTransformer. Remarkably, the HSQformer matched the diagnostic accuracy of senior radiologists and comprehensively outperformed that of junior radiologists. These findings underscore the robustness and clinical relevance of our AI-assisted diagnostic tool in HCC screening.
@@
  • To the best of our knowledge, this study is the first systematic investigation into the application of AI-assisted ultrasound for HCC screening. We have conducted a comprehensive retrospective data analysis from 2014 to 2022 across three clinical scenarios—single-center, multi-center, and high-risk patient testing—to meticulously assess the AI tool’s efficacy and clinical applicability.
@@
  • We have established the pioneering benchmark for AI-assisted ultrasound screening of HCC, which is accessible at https://github.com/Asunatan/HSQformer.
@@ 2.1 Convolutional Neural Networks
The advent of CNNs has been pivotal in the field of computer vision, establishing their architecture as a cornerstone for critical tasks such as image classification, object detection, and semantic segmentation[25, 26]. CNNs’ widespread adoption stems from their ability to automatically learn complex features and patterns directly from raw images, thereby eliminating the reliance on manual feature engineering. Inherent inductive biases, notably locality and translational invariance, are crucial for effectively addressing a broad spectrum of visual recognition tasks. These biases enable CNNs to concentrate on local feature patterns within images, allowing the network to capture critical visual information while maintaining robustness against positional or scale variations. However, despite their success, CNNs have certain limitations, particularly in the domain of medical imaging. One significant drawback is their restricted receptive field, which limits their ability to capture global semantic information—a critical aspect in medical image analysis where comprehensive understanding of the entire image context is essential for accurate diagnosis. To enhance the generalizability of CNNs, several pioneering approaches have been introduced, such as dilated convolutions (Atrous convolutions)[27], deformable convolutions[11, 12], and attention mechanisms like SENet[28] and CBAM[29], which recalibrate channel-wise features and enhance both channel and spatial attention, respectively.
@@ 2.2 Vision Transformers
In natural language processing (NLP), Transformers[30] have demonstrated exceptional capabilities in modeling long-range dependencies. This success has inspired researchers to explore the application of transformer architectures to computer vision, where the ViT represents a landmark development. Recent works have explored applying ViT to various vision tasks: image classification, object detection, image segmentation, depth estimation, image generation, video processing, and others. Despite its success, ViT also presents certain limitations. One notable drawback is its reliance on large-scale datasets for effective training, as Transformers lack the inductive biases inherent in CNNs, such as locality and translation invariance[31]. As a result, ViTs may struggle to generalize effectively when trained on smaller datasets[32]. Additionally, ViTs are computationally demanding, particularly due to their self-attention mechanisms, which scale quadratically with image size, leading to high memory and processing costs. To enhance the efficiency of ViTs while maintaining performance, several studies[16, 17, 18, 19, 20, 21] have proposed methods that incorporate local self-attention mechanisms or pooling operations. By restricting self-attention to localized regions of an image or employing pooling to condense feature maps, these optimizations not only streamline processing but also ensure that the models remain adept at handling diverse vision tasks, thereby achieving a balance between computational efficiency and model efficacy. Several other studies have primarily concentrated on diversifying the ViT architectures and enhancing their adaptability, such as the Focal Transformer[33], MobileViT[34], and Axial-Attention Transformer[35], each contributing to the field with innovative approaches to attention mechanisms and knowledge distillation techniques.
@@ 2.3 Latent Space and Sparse Learning
Latent Space is central to capturing intrinsic data features in machine learning, with its utility in deep learning models demonstrated through intermediate layer representations that are crucial for various tasks[36]. Learned queries, as explored in SetTransformers[37] and Perceiver [38, 39] networks, project inputs into a lower-dimensional space for efficiency and prediction accuracy. Goyal et al.[36] enhance transformers by using learned queries as a shared workspace, reducing computational complexity. Self-attention mechanisms have also been optimized with learnable tokens replacing queries or keys, as shown in Involution[40], VOLO[41], and QnA[42], to generate dynamic affinity matrices and improve model performance. These innovations underscore the importance of Latent Space in advancing deep learning capabilities.
Sparse learning stands as a pivotal approach within machine learning, particularly for its effectiveness in optimizing computational resources and enhancing model performance. A prime exemplar of this is the Mixture-of-Experts (MOE) framework, which segments models into specialized experts, selectively activating only the relevant subset to process inputs, thereby achieving computational efficiency. The Multi-Mixture-of-Experts (MMoE)[43] model epitomizes MOE’s application in multi-task learning, designed to explicitly learn task relationships from data. It enables the sharing of expert sub-models across various tasks while simultaneously training a gating network to optimize performance for each individual task. Within large-scale models, MOE has attracted considerable scholarly attention; studies such as Qwen-MOE[44], DeepSeek[45, 46], and LLaMA-MoE[47, 48] are at the forefront of research, showcasing the potential of MOE in scaling up large language models (LLMs) with a consistent number of activated parameters. In summary, MOE emerges as a potent sparse representation model, and it is recommended that readers consult reference [49] for an in-depth exploration of the subject.
@@ 3 Approach
Figure 2: Overview of the HSQformer, integrating CNN and ViT features through a hierarchical sparse querying framework for efficient diagnosis.
@@ 3.1 Overall Architecture
Our objective is to design an architecture that effectively combines the complementary strengths of CNNs and ViTs, devoid of any unnecessary bells-and-whistles. This architectural framework is deployed in real-world clinical ultrasound screening for hepatocellular carcinoma (HCC) to investigate the potential of AI-assisted diagnosis in a clinical context. Previous studies[23] have indicated that combining CNNs and ViTs may lead to feature redundancy. To address this issue, we introduce a paradigm of latent space representation and sparse learning. As illustrated in Figure 2, our architecture consists primarily of a feature extractor, a projector, and a HSQformer. The HSQformer is a novel multi-level sparse Q-former, composed of stacked hierarchical learnable Query Transformers. We selected the state-of-the-art ConvNeXt and SwinTransformer as feature extractors due to their simplicity and efficiency. Similar to ConvNeXt and Swin Transformer, the HSQformer is organized into four stages that aggregate feature maps at various scales. Each stage shares a similar modular design, consisting of multiple Cross-Self-attention Mixed experts (CSM) blocks, which enables scalability and reusability.
@@ 3.2 HSQformer
To harness the complementary strengths of ConvNeXt and SwinTransformer, a straightforward approach is to simply sum their features. However, this method suffers from feature redundancy, and our preliminary experiments, as depicted in Figure 1, show that the benefits are not significant. To address these issues, inspired by references [42, 50, 43], we introduce latent space representation and sparse learning to construct a novel hierarchical sparse querying transformer, termed HSQformer. This architecture comprises a learnable query embedding representation and a four-stage backbone, with each stage composed of Cross-Self-attention Mixed experts (CSM) Modules, facilitating the scalability of the HSQformer. The size of the HSQformer can be varied by adjusting the number of CSM modules in each stage, resulting in different configurations such as Small, Base, and Large with detailed specifications provided in Appendix.
@@ 3.2.1 CSM
As illustrated in Figure 2, the Cross-Self-attention Mixed experts (CSM) consists of cross-attention, self-attention, and a mixture of experts, closely aligning with the standard transformer architecture without incorporating unnecessary bells and whistles. The role of CSM varies depending on its specific location within the model. When situated within the projector, the CSM facilitates the integration of fine-grained local features with global contextual information, and projects the resulting representations into a $D$-dimensional latent space for subsequent processing. In contrast, when integrated into the HSQformer, the CSM primarily leverages query embeddings to extract features from the latent space that are highly pertinent to diagnostic classification tasks, ensuring that the most relevant information is captured for accurate decision-making.
@@ 4.1 Experimental Setup
The dataset utilized in this study was sourced from the multi-center dataset of the Ultrasound Engineering Society of China’s Medical Industry Branch (UE-MICAP), which includes 19,464 images collected from January 2014 to December 2022. The data were obtained from the following hospitals: the First Affiliated Hospital of Sun Yat-sen University, the Third Affiliated Hospital of Sun Yat-sen University, the Sixth Affiliated Hospital of Sun Yat-sen University, the Seventh Affiliated Hospital of Sun Yat-sen University, the First Affiliated Hospital of Guangzhou Medical University, the First Affiliated Hospital of Guangxi Medical University, Foshan Sanshui District People’s Hospital, and West China Xiamen Hospital of Sichuan University. The inclusion criteria for the study were: (a) patients at risk of HCC, including those with clinical diagnoses or imaging evidence of cirrhosis, or pathological confirmation of cirrhosis; (b) males over the age of 40 and females over the age of 50 with a history of viral hepatitis; (c) patients who underwent ultrasound screening and had lesions confirmed by clinical or pathological diagnosis. The exclusion criteria were: (a) individuals under the age of 18; (b) images with inadequate quality; (c) images containing both benign and malignant lesions, as this could potentially impair the AI’s ability to accurately classify the lesions. This retrospective study utilized fully anonymized images, thus eliminating the need for informed consent.
Table 1 outlines the three testing scenarios established in this study: Single-Center Test, Multi-Center Test, and High-Risk Patient Test. The Single-Center Test refers to an independent external center evaluation using a dataset distinct from the training set. For the Multi-Center Test, data were aggregated from eight different hospitals to simulate a more diverse and generalized clinical environment. The High-Risk Patient Test specifically focuses on patients with a high risk of hepatitis, aiming to assess model performance within a particularly challenging and clinically relevant subgroup. These testing scenarios have been designed to closely mirror real-world clinical applications, thereby providing a comprehensive assessment of the model’s effectiveness and robustness across varied conditions.
@@
Benign 8225 200 2295 706
Malignant 3155 435 2270 2178
@@
Table 2: This table presents a comprehensive comparison of our proposed model with existing state-of-the-art CNNs and ViT models, highlighting their respective performances and characteristics.
@@
The original DICOM images were losslessly converted to PNG format using OpenCV, standardizing the image format for subsequent processing. Subsequently, textual metadata, including patient and device identifiers, institutional information, and date stamps, were removed from the images utilizing the YOLO object detection model[55]. This study leveraged the PyTorch 2.3.1 deep learning framework, with computations accelerated on an NVIDIA A100 GPU for both training and evaluation purposes. To ensure a fair comparison, we followed the fine-tuning protocol as detailed in CSWinTransformer[18]. Specifically, all models underwent pre-training on the ImageNet dataset, with an input resolution set at 224×244 pixels. Following pre-training, fine-tuning was conducted using the AdamW optimizer, employing a weight decay of 1e-4 over 32,000 iterations, while maintaining the aforementioned input resolution. The default settings for batch size and initial learning rate were 18 and 1e-4, respectively, with the learning rate adjusted according to a cosine annealing schedule. To enhance dataset diversity and mitigate overfitting, data augmentation techniques such as horizontal flipping, ColorJitter, and Mixup were applied during training. A 5-fold patient-level cross-validation strategy was implemented to prevent any data leakage between the training and validation sets, ensuring robustness and generalizability of the results. Model weights were saved after each epoch if there was an improvement in validation performance, measured by the area under the receiver operating characteristic curve (AUC). Notably, early stopping was not employed, to avoid potential underfitting or overfitting of the models to the dataset. The appendix provides additional detailed information.

### v2

@@ Abstract
Hepatocellular carcinoma (HCC), ranking as the third leading cause of cancer-related mortality worldwide, demands urgent improvements in early detection to enhance patient survival. While ultrasound remains the preferred screening modality due to its cost-effectiveness and real-time capabilities, its sensitivity (59%-78%) heavily relies on radiologists’ expertise, leading to inconsistent diagnostic outcomes and operational inefficiencies. Recent advancements in AI technology offer promising solutions to bridge this gap. This study introduces the Hierarchical Sparse Query Transformer (HSQformer), a novel hybrid architecture that synergizes CNNs’ local feature extraction with Vision Transformers’ global contextual awareness through latent space representation and sparse learning. By dynamically activating task-specific experts via a Mixture-of-Experts (MoE) framework, HSQformer achieves hierarchical feature integration without structural redundancy. Evaluated across three clinical scenarios—single-center, multi-center, and high-risk patient cohorts—HSQformer outperforms state-of-the-art models (e.g., 95.38% AUC in multi-center testing) and matches senior radiologists’ diagnostic accuracy while significantly surpassing junior counterparts. These results highlight the potential of AI-assisted tools to standardize HCC screening, reduce dependency on human expertise, and improve early diagnosis rates. The full code is available at https://github.com/Asunatan/HSQformer.
footnotetext: * Equal contribution.footnotetext: 🖂Corresponding Author.
@@ 1 Introduction
Liver cancer ranks as the sixth most common malignancy and the third leading cause of cancer-related mortality globally, with hepatocellular carcinoma (HCC) accounting for approximately 75% - 85% of primary liver cancers[1]. Early detection and treatment of HCC can improve patient survival rates and life expectancy. Ultrasound (B-mode) is the preferred imaging modality for screening high-risk populations for HCC. Compared to CT or MRI, ultrasound offers the advantages of portability, radiation-free imaging, and real-time visualization. However, according to meta-analyses, the sensitivity of conventional B-mode ultrasound in detecting HCC ranges from only 59% to 78%, which is highly dependent on the clinical experience of radiologists[2, 3]. Moreover, ultrasound examinations require substantial time from these specialists for image interpretation and secondary confirmation, often making clinical assessments inefficient. These limitations underscore the urgent need for robust tools to standardize HCC screening and reduce reliance on subjective human interpretation.
Recent advancements in artificial intelligence (AI), particularly deep learning-based computer-aided diagnosis (CAD) systems, demonstrate transformative potential for improving clinical decision-making in medical image interpretation.
@@
Current AI-based CAD frameworks predominantly leverage convolutional neural networks (CNNs) or vision transformers (ViTs)[4]. CNNs have been pivotal in computer vision tasks due to their ability to capture local texture information, but their performance is often limited by the size of their receptive fields, restricting their ability to perceive long-range dependencies and integrate global contextual information. ViTs, on the other hand, excel in capturing global semantic information through self-attention mechanisms but typically require extensive datasets and significant computational resources for effective pre-training, and lack the inductive biases inherent in CNNs.
@@
This raises the question: is there a straightforward method to address the limitations of these two architectures? A simple yet feasible solution is to combine the strengths of CNNs and ViTs to form a hybrid model. However, current integration methods often necessitate complex architectural modifications or the use of sophisticated design techniques. Is it possible to achieve effective complementarity between CNNs and ViTs without altering the network structure? This task presents significant challenges. As shown in Figure 1, the two most popular network architectures, ConvNext (CNNs)[5] and SwinTransformer (ViTs)[6], were selected for preliminary experiments. Our results show that late fusion[7] of CNNs and ViTs offers limited benefits, as this approach tends to overemphasize high-level semantics for classification while neglecting low-level details and textures. Moreover, this approach[7] fails to leverage multi-scale features, which have been proven crucial for downstream tasks such as segmentation and detection.
To address these limitations and bridge the gap between the two architectures, we propose the Hierarchical Sparse Query Transformer (HSQformer), a novel hybrid architecture that hierarchically synergizes the local feature extraction capabilities of CNNs with the global contextual awareness of ViTs through latent space representation and sparse learning. The HSQformer dynamically activates task-specific experts via a Mixture-of-Experts (MoE) framework, achieving hierarchical feature integration without structural redundancy. Evaluated across three clinical scenarios—single-center, multi-center, and high-risk patient cohorts—the HSQformer demonstrates superior performance compared to state-of-the-art models and matches the diagnostic accuracy of senior radiologists while significantly surpassing that of junior counterparts. This study represents the first systematic exploration of AI-assisted ultrasound screening for HCC, offering a robust and standardized diagnostic tool to improve early detection rates and patient outcomes.
@@
  • To our best knowledge, our work represents the first large - scale validation of AI - assisted ultrasound screening for HCC across distinct clinical environments, ensuring generalizability and practical relevance.
@@
  • We have open-sourced our code and checkpoints, and established the pioneering benchmark for AI-assisted ultrasound screening of HCC, which is accessible at https://paperswithcode.com/sota/classification-on-liver-us.
@@ 2.1 Convolutional Neural Networks
Convolutional Neural Networks (CNNs) have revolutionized computer vision, serving as a cornerstone for tasks such as image classification, object detection, and semantic segmentation [8, 9]. Their success stems from their ability to autonomously learn hierarchical features directly from raw pixel data, eliminating the need for manual feature engineering. Key inductive biases, including spatial locality and translational invariance, enable CNNs to efficiently capture local patterns while maintaining robustness to positional variations in natural images.
Despite these strengths, conventional CNNs exhibit limitations in medical imaging applications. Their restricted receptive fields hinder the integration of global contextual information, which is critical for interpreting complex anatomical structures and subtle pathological features [3]. This constraint becomes particularly detrimental when local regions require semantic dependencies spanning the entire image.Recent advancements aim to enhance CNN generalizability through three synergistic strategies: (1) Expanding receptive fields via enlarged kernels [10, 11, 12] or structural re-parameterization[13, 14, 15]; (2) Incorporating adaptive operators such as dilated[16] or deformable convolutions [17, 18, 19] to dynamically adjust spatial sampling; (3) Integrating attention mechanisms[20, 21] to refine feature discriminability. These innovations preserve CNNs’ local feature extraction strengths while enhancing global contextual modeling, positioning them as adaptable frameworks for medical image analysis
@@ 2.2 Vision Transformers
In natural language processing (NLP), Transformers[22] have demonstrated exceptional capabilities in modeling long-range dependencies. This success has inspired researchers to explore the application of Transformer architectures to computer vision, where the ViT represents a landmark development. Recent works have explored applying ViT to various vision tasks: image classification, object detection, image segmentation, depth estimation, image generation, video processing, and others. Despite its success, ViT also presents certain limitations. One notable drawback is its reliance on large-scale datasets for effective training, as Transformers lack the inductive biases inherent in CNNs, such as locality and translation invariance[23]. As a result, ViTs may struggle to generalize effectively when trained on smaller datasets[24]. Additionally, ViTs are computationally demanding, particularly due to their self-attention mechanisms, which scale quadratically with image size, leading to high memory and processing costs.
To enhance the efficiency of ViTs while maintaining performance, several studies[6, 25, 26, 27, 28, 29] have proposed methods that incorporate local self-attention mechanisms or pooling operations. By restricting self-attention to localized regions of an image or employing pooling to condense feature maps, these optimizations not only streamline processing but also ensure that the models remain adept at handling diverse vision tasks, thereby achieving a balance between computational efficiency and model efficacy. Several other studies have primarily concentrated on diversifying the ViT architectures and enhancing their adaptability, such as the Focal Transformer[30], MobileViT[31], and Axial-Attention Transformer[32], each contributing to the field with innovative approaches to attention mechanisms and knowledge distillation techniques.
@@ 2.3 Latent Space and Sparse Learning
The concept of Latent Space is pivotal in machine learning for capturing the intrinsic features of data. Its utility in deep learning models is exemplified by intermediate layer representations, which are crucial for various tasks[33]. To enhance efficiency and prediction accuracy, learned queries have been explored in architectures such as SetTransformers[34] and Perceiver [35] networks, where they project inputs into a lower-dimensional space. Goyal et al.[33] further advanced this concept by leveraging learned queries as a shared workspace to reduce computational complexity in Transformers. In addition to these developments, attention mechanisms have also been optimized through the incorporation of learnable tokens. Models such as Involution[36], VOLO[37], and QnA[38] demonstrate how learnable tokens can replace traditional queries or keys, generating dynamic affinity matrices that significantly boost model performance. Collectively, these innovations underscore the importance of Latent Space in driving the evolution and enhancing the capabilities of deep learning models.
Sparse learning stands as a pivotal approach within machine learning, particularly for its effectiveness in optimizing computational resources and enhancing model performance. A prime example of this approach is the Mixture-of-Experts (MOE) framework, which partitions models into specialized experts, selectively activating only the relevant subset to process inputs and thereby achieving computational efficiency. The Multi-Mixture-of-Experts (MMoE)[39] model epitomizes MOE’s application in multi-task learning, designed to explicitly learn task relationships from data. It enables the sharing of expert sub-models across various tasks while simultaneously training a gating network to optimize performance for each individual task. Within the realm of large language model(LLM), MOE has attracted considerable scholarly attention; studies such as DeepSeek[40], and LLaMA-MoE[41, 42] are at the forefront of research, showcasing the potential of MOE in scaling up large language models with a consistent number of activated parameters. In summary, MOE emerges as a potent sparse representation model, and it is recommended that readers consult reference [43] for an in-depth exploration of the subject.
@@ 3 Approach
Figure 2: model architecture. Overview of the HSQformer, integrating CNN and ViT features through a hierarchical sparse querying framework for efficient diagnosis.
@@ 3.1 Overall Architecture
Our objective is to design an architecture that effectively combines the complementary strengths of CNNs and ViTs, without any unnecessary bells-and-whistles design. This architectural framework is deployed in real-world clinical ultrasound screening for hepatocellular carcinoma (HCC) to investigate the potential of AI-assisted diagnosis in a clinical context. To achieve this goal, we introduce a paradigm that leverages latent space representation and sparse learning to integrate these strengths. As illustrated in Figure 2, our architecture consists primarily of a feature extractor, a projector, and a HSQformer. The HSQformer is a novel multi-level sparse Q-former, composed of stacked hierarchical learnable Query Transformers. We selected the state-of-the-art ConvNeXt and SwinTransformer as feature extractors due to their simplicity and efficiency. Similar to ConvNeXt and SwinTransformer, the HSQformer is organized into four stages that aggregate feature maps at various scales. Each stage shares a similar modular design, consisting of multiple Cross-Self-attention Mixed experts (CSM) blocks, which enables scalability and reusability.
@@ 3.2 HSQformer
Inspired by references [38, 44, 39], we introduce latent space representation and sparse learning to construct a novel Hierarchical Sparse Querying Transformer, termed HSQformer. This architecture comprises a learnable query embedding representation and a four-stage backbone, with each stage composed of Cross-Self-attention Mixed experts (CSM) modules, facilitating the scalability of the HSQformer. The size of the HSQformer can be varied by adjusting the number of CSM modules in each stage, resulting in different configurations such as Small, Base, and Large with detailed specifications provided in Appendix.
@@ 3.2.1 CSM
As illustrated in Figure 2, the Cross-Self-attention Mixed experts (CSM) consists of cross-attention, self-attention, and a mixture of experts, closely aligning with the standard transformer architecture without incorporating unnecessary bells-and-whistles. The role of CSM varies depending on its specific location within the model. When situated within the projector, the CSM facilitates the integration of fine-grained local features with global contextual information, and projects the resulting representations into a $D$-dimensional latent space for subsequent processing. In contrast, when integrated into the HSQformer, the CSM primarily leverages query embeddings to extract features from the latent space that are highly pertinent to diagnostic classification tasks, ensuring that the most relevant information is captured for accurate decision-making.
@@ 4.1 Experimental Setup
The dataset in this study was sourced from the multi-center dataset of the Ultrasound Engineering Society of China’s Medical Industry Branch (UE-MICAP). It contains 11,149 cases with 19,464 images collected from January 2014 to December 2023. The data were obtained from these hospitals: the First, Third, Sixth, and Seventh Affiliated Hospitals of Sun Yat-sen University; the First Affiliated Hospital of Guangzhou Medical University; the First Affiliated Hospital of Guangxi Medical University; Foshan Sanshui District People’s Hospital; and West China Xiamen Hospital of Sichuan University.
The inclusion criteria for the study were: (a) patients at risk of HCC, including those with clinical diagnoses or imaging evidence of cirrhosis, or pathological confirmation of cirrhosis; (b) males over the age of 40 and females over the age of 50 with a history of viral hepatitis; (c) patients who underwent ultrasound screening and had lesions confirmed by clinical or pathological diagnosis. The exclusion criteria were: (a) individuals under the age of 18; (b) images with inadequate quality; (c) images containing both benign and malignant lesions, as this could potentially impair the AI’s ability to accurately classify the lesions. This retrospective study utilized fully anonymized images, thus eliminating the need for informed consent.
Table 1 presents the three testing scenarios: single-center test, multi-center test, and high-risk patient test. The single-center test is an independent evaluation using a dataset separate from the training set. For the multi-center test, data from eight hospitals were combined to mimic a diverse clinical setting. The high-risk patient test focuses on patients with hepatitis, assessing model performance in this challenging subgroup. These scenarios mirror real-world clinical applications, offering a comprehensive evaluation of the model’s effectiveness and robustness across different conditions.
@@
Benign images 8225 200 2295 706
Malignant images 3155 435 2270 2178
@@
Table 2: SOTA Model Comparison. This table presents a comprehensive comparison of our proposed model with existing state-of-the-art CNNs and ViT models, highlighting their respective performances and characteristics.
@@
The original DICOM images were losslessly converted to PNG format via OpenCV for standardization. Then, textual metadata like patient identifiers, device information, institutional details, and date stamps were removed using the YOLO object detection model[50]. We followed the fine-tuning protocol from CSWinTransformer[26] for fair comparison. A 5-fold patient-level cross-validation strategy was implemented to prevent data leakage and ensure result robustness. Model weights were saved after each epoch with AUC - based validation performance improvement, without early stopping to avoid underfitting or overfitting. More details are in the appendix.